OpenStudy (nice):

how to find the projection of a=(1,3) on b=(4,0,2) ??

OpenStudy (amistre64):

a onto b eh... vector right? its the |a| cos(angle between)

OpenStudy (nice):

really ? I know diff. formula!!

OpenStudy (nice):

angle is not mentioned at all in the question

OpenStudy (amistre64):

t = \(cos^{-1}(a.b/(|a||b|))\)

OpenStudy (amistre64):

well, in order to know the length of the projection onto b you gotta know the angle between a and b

OpenStudy (nice):

not the length ! it said just the vector that is proj.

OpenStudy (amistre64):

vectors are length and direction...cant have vector without length right?

OpenStudy (nice):

sure! but I'm using the dot product ,, (a.b/||b||^2) x b

OpenStudy (amistre64):

ok...

OpenStudy (amistre64):

<1,3,0> <4,0,2> ------- 4+0+0 =4

OpenStudy (nice):

aha ,, then ?

OpenStudy (amistre64):

sqrt(10).sqt(20) = sqrt(200)

OpenStudy (amistre64):

cos_1(4/sqrt(200)) is what id try

OpenStudy (amistre64):

73.57 degrees is the angle between them...

OpenStudy (amistre64):

but we already have the cos of that: 4/sqrt(200) lol

OpenStudy (amistre64):

4sqrt(10) -------------- = 4/sqrt(20) right? this is the length of a vector sqrt(10).sqrt(20) that is parallel to b

OpenStudy (amistre64):

b/|b| = unit vector of b; then *that to get the vector projection

OpenStudy (nice):

aha that the answer I've found but in diff. way

OpenStudy (nice):

but the problem that it's multiple choice Qs. and I didn't find the answer !!

OpenStudy (amistre64):

<4,0,2> 4 ------- * ----- :) sqrt(20) sqrt20

OpenStudy (amistre64):

whats the choices?

OpenStudy (amistre64):

i get: <16/20 , 0 ,8/20> if im right lol

OpenStudy (nice):

the answers are : <1,1> , <2,1> , <2,-1> , <1,2>

OpenStudy (nice):

that's the answer I've got ! but it seems that its wrong !!!

OpenStudy (amistre64):

well, then we misunderstood what it was asking for then :)

OpenStudy (nice):

the question which is up there is the exact same ques. !!!!!

OpenStudy (amistre64):

yeah, perhaps its the vector that I drew in green?

OpenStudy (nice):

thanks allot for trying ..

OpenStudy (amistre64):

a.b --- * b |b|^2 4/20 = 1/5<4,0,2> = <4/5,0,2/5> cant really see it being nuthin else

OpenStudy (amistre64):

youre welcome :) good luck

OpenStudy (nice):

yes! that's my opinion two !

OpenStudy (nice):

thank you ..